Hall A CC Meeting Minutes - May 18, 2000

Present: Kees de Jager, Ron Ransome, Ed Brash, Christian Cavata, Franco Garibaldi

1. Status Report - Kees de Jager

The 3He(e,e'p) + parasitic 3He(e,e'pn) experiments have been completed.
The 4He(e,e'p) polarization transfer experiment (E93049) has been completed.

Currently, preparations are underway for the N-Delta experiment, which will
run until the August shutdown.

a) The cooling capacity for the 3He target is much less than expected.  The 
solution is to run two heat exchangers in parallel.  This allows to run beam
currents in excess of 100 uA.  However, this is not a permanent solution.  
There is a new heat exchanger design; the plan is to install this in 
September so as to be ready for Templon's experiment.

b) There is a problem with positioning due to an upstream girder which is 
fixed to newly installed quadrupole magnets.  Unfixing the girder resulted
in significant movement.  Currently working on other possible fixes.

c) There was a small leak in the H-target beer can.  This was fixed.

Plans for the Summer Shutdown
a) Switch detector stacks (for Gep experiment).
b) Install new heat exchanger.
c) Install a new upstream girder for the septa (in advance).

There is a general long term problem with the alignment of the spectrometers
and the beamline.  This is currently being investigated.  Currently, the 
girder position is not stable against bumps from work in the hall (jumps on
the order of 0.8 mm, require 0.1 mm for many experiments).

The energy measurements disagree at about 1.5 x 10^-3 level (Nilanga's 
energy measurement from spectrometers is close to e-p measurement.

Also, Q1 position is unstable when ladders are placed against it for other 

2. New Members

Riad Suleiman, Doug Higinbotham, and Alan Nathan's post-doc have been put
forward.  They are to be voted on at the collaboration meeting.

3.  DVCS proposal

Generally, all of the reviews were favourable.  There were some concerns about
rates in the photon calorimeter.  Also, some time for test data may be useful.

--> The proposal was unanimously supported by the CC as a Hall A Collaboration

4.  CC Elections

The new members of the CC are Adam Sarty and Bob Michaels.

5.  PAC 19

The CC expects 5-6 resubmissions.  The deadline for Hall A submissions is 
expected to be around mid-November.

6.  Plan for Progress Report

The next progress report (deadline for submissions sometime in September) 
should concentrate on new items which were not included in last year's 
report.  In addition, there will be reports from all experiments which have
completed data taking.  It was decided NOT to post the report on the web
at this time.

7.  Referee Reports on Papers

It was unanimously recommended by the CC that any reports from referees on
submitted papers should be made available to ALL AUTHORS by the spokespersons
of the experiment as soon as possible after they are received, and certainly
prior to a response being submitted.